POGI Company sells television at an average price of P7, 500 and also offers to each customer a separate 3-year warranty contract for P750 that requires the company to perform periodic services and to replace defective parts. During 2020, the company sold 300 television and 270 warranty contracts for cash. It estimates the 3-year warranty costs as P200 for parts and P400 for labor and
accounts for warranties separately, assumes sales occurred on December 31, 2020, income is
recognized on the warranties, and straight line recognition of warranty revenues occurs.
a. What amount of current liability relative to warranty revenue would appear on the December 31,
2020 balance sheet, respectively?
b. What amount of non-current liability relative to warranty revenue would appear on the December
31, 2020 balance sheet, respectively?

Lastly, Hudson Hotel collects 15% in city sales taxes on room rentals, in addition to a Php 2 per room, per night, occupancy tax. Sales taxes for each month are due at the end of the following month, and occupancy taxes are due 15 days after the end of each calendar quarter. On January 3, 2021, Hudson paid its November 2020 sales taxes and its fourth quarter 20x0 occupancy taxes. Additional information pertaining to Hudson's operations is:
2020            Room rentals          Room nights
October          100,000                    1,100
November      110,000                     1,200
December       150,000                    1,800
a. What amount should Hudson report as sales taxes payable?
b. What amount should Hudson report as Occupancy taxes payable in its December 31, 2020, balance sheet?
Answer #1

POGI Company

A) The amount of current liability relative to warranty revenue that would appear on the December 31,
2020 balance sheet is

Current liabilities :

Unearned warranty revenue $67500

B) The  amount of non current liability relative to warranty revenue that would appear on the December 31,
2020 balance sheet is

Non current liability

Unearned warranty revenue $ 135000

Workings

Unearned warranty revenue = 270 * 750 = 202500

Current liability ( 3 year Warranty ) = 202500 / 3 = 67500

Non current liability ( 3 year warranty ) = (202500*2)/3 = 135000